Discuss with Askiitians Tutors

HOW MUCH TOUGH IS NSEP,NSEC,RMO WITH RESPECT TO BOARD(TAKE IT AS 2/10 IN DIFFICULTI SCALE) AND JEE MAINS(6/10 IN DIFFICULTI SCALE).PLS GIVE THE ANSWER SEPARATELY FOR THE SEPARATE EXAMS MENTIONED ABOVE???

Avirup Kumar Roy
10 Years agoGrade

1 Answer

Sourabh Singh
10 Years ago
NSEP = 5/10

NSEC = 5/10

RMO = 5/10